OpenStudy (anonymous):

Can someone help me figure out this? What is the domain of validity of the trigonometric identity csc theta = 1 / sin theta?

OpenStudy (anonymous):

@UnkleRhaukus @MathLegend

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

I know it is a reciprocal identity I just cannot figure out how to find the domain of validity

hero (hero):

Well, we know sin theta cannot equal zero. So find the angles that would make sin theta equal zero.

OpenStudy (anonymous):

when I type it into my calculator, when x or theta =0 I get an error ?

hero (hero):

What did you type into your calculator?

OpenStudy (anonymous):

I typed 1/sin (theta)

OpenStudy (anonymous):

into a graphing calculator, I looked at the table and it said error for x being 0

hero (hero):

Okay, so you know that zero is one of the angles that will not work.

OpenStudy (anonymous):

Thats what my graph looks like, it doesn't touch zero at all ?

hero (hero):

Do this: Try graphing 1/sin(theta) in one color. Then graph csc(theta) in another color. See if you can notice a distinction between the two graphs.

OpenStudy (anonymous):

My calculator doesn't have a csc button,

hero (hero):

What kind of calculator do you have?

hero (hero):

I'm pretty sure that if you're using a graphing calc, it has one.

OpenStudy (anonymous):

TI-84 graphing callculator.

OpenStudy (anonymous):

I don't see it anywhere.

hero (hero):

Yup, it has one.

OpenStudy (anonymous):

Where?! Our teachers been telling us to use the 1/ equation.

hero (hero):

Yes, because there is no distinction between csc(theta) and 1/sin(theta)

OpenStudy (anonymous):

I understand that.

hero (hero):

By the way, I suggest you get a graphing calc with a csc function on it.

OpenStudy (anonymous):

They don't give us one for my online school.

hero (hero):

Do you go to connections academy?

OpenStudy (anonymous):

yes

hero (hero):

I see. Seems like everyone does.

OpenStudy (anonymous):

I know!!!

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@kymber

OpenStudy (phi):

What is the domain of validity of the trigonometric identity csc(x) The domain are all values of x that give a defined value for csc(x) (1/sin(x)) we run into trouble if we take the square root of a negative number (not an issue here) or divide by zero. Dividing by 0 is a problem here: any time sin(x) is 0, 1/sin(x) is a divide by zero. so we must exclude x=0º because sin(0)=0 and 1/sin(0) is not defined.

OpenStudy (phi):

However, sin(x) is 0 a more than just sin(0), At every 180 degrees (or pi radians) sin is 0

OpenStudy (anonymous):

so I've narrowed it down to either a.the set of real numbers except for zero and odd multiples of pi/2 or b.the set of real numbers except for zero and odd multiples of pi

OpenStudy (anonymous):

am I on the right track?

OpenStudy (phi):

In radians, we would exclude, ... -3π, -2π, -π, 0, π, 2π, 3π ...

OpenStudy (anonymous):

so it would be b?

OpenStudy (phi):

you should know that pi/2 is ok, because sin(pi/2) = 1 so you are not dividing by 0

OpenStudy (phi):

Isn't there an option for all multiples of pi?

OpenStudy (phi):

unless they are being tricky: an even multiple of pi/2 is 2*pi/2 or 4*pi/2 which simplifies to pi, 2pi, etc

hero (hero):

I tried telling her to find all the angles that would make sin theta equal zero. I just didn't want to give her the answer.

OpenStudy (phi):

It is a fine line between helping just enough and too much...

OpenStudy (anonymous):

I just thought that 0 would not be included in the domain? along with all the multiples of pi?

hero (hero):

How come you didn't say that earlier when I was helping you?

OpenStudy (phi):

yes, but what choice is that? the 2 you listed don't match that

OpenStudy (anonymous):

the set of real numbers the set of real numbers except for multiples of pi the set of real numbers except for zero and odd multiples of pi the set of real numbers except for zero and odd multiples of pi/2

OpenStudy (phi):

They are being just a little tricky. we can assume that multiples of pi include 0 (we have to, none of the other choices are correct) odd multiples of pi pi,3pi,5pi... leave out 2pi,4pi, ... which would cause a divide by 0

OpenStudy (anonymous):

so it would be the 2nd choice.?

OpenStudy (phi):

Yes, do you see why?

hero (hero):

zero is a multiple of pi because 0*pi = 0

OpenStudy (anonymous):

Because even the even multiples cause an invalidity in the equation, not just the odd multiples of pi.

OpenStudy (phi):

and what causes the invalidity?

OpenStudy (anonymous):

thats what I don't understand about this entire unit, I've asked my teacher multiple times and she keeps telling me to reread the section.

OpenStudy (phi):

It is not too esoteric: dividing by a 0 is not allowed (people don't know what it means) so the invalid expression would happen if you divide by 0.

hero (hero):

Cause of invalidity: \[\csc(\pi) = \frac{1}{\sin(180^{\circ})} = \frac{1}{0}\]

OpenStudy (anonymous):

So, that mean that any number that able to be divided by 0?

OpenStudy (phi):

if we defined a function f(x)= 1/x this function will not be defined at x=0 (divide by 0 not allowed). Its domain would exclude x=0 same idea for 1/sin(x) except sin(x) is zero at lots of places.. k*pi , k an integer

OpenStudy (anonymous):

oh! That diagram actually helps a little.

hero (hero):

pi radians = 180 degrees \[\csc(\pi) = \frac{1}{\sin(\pi)} =\frac{1}{\sin(180^{\circ})} = \frac{1}{0}\] That should make it a little more clearer

OpenStudy (phi):

So, that mean that any number that able to be divided by 0? can you ask again?

hero (hero):

so If pi makes the expression invalid, then ANY multiple of pi will also, including zero.

OpenStudy (anonymous):

Okay, so no matter what which expression of pi, odd or even, would make the expression invalid?

hero (hero):

We know pi is odd zero degrees equals 2 pi which is even.

OpenStudy (phi):

Okay, so no matter what which expression of pi, odd or even, would make the expression invalid? to be clear, I would say makes sin(x) 0 and 1/sin(x) would be 1/0 which is invalid

OpenStudy (anonymous):

Thank you, my teacher has not taken any time to help me out and it is really starting to irritate me. I'm two weeks behind in this class because I don't understand it and everytime I ask for clarification she tells me to reread the chapter.